# Finely Stratified Rerandomization Designs

### v1

@@ 5.1 Minimax Rerandomization
Suppose that we know $\gamma_{0}\in B$ for some prior information set $B\subseteq\mathbb{R}^{d_{h}}$. Fix $\epsilon>0$ and consider a “minimax” style acceptance criterion, rerandomizing the treatments $D_{1:n}$ until
$\sup_{\gamma\in B}|\gamma^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]|\leq\epsilon.$ (5.1)
Note that the function $f_{B}(x)=\sup_{\gamma\in B}|\gamma^{\prime}x|$ is convex, so we can also interpret this as a convex imbalance penalty, rerandomizing until $f_{B}(\mathcal{I}_{n})\leq\epsilon$ for imbalance metric $\mathcal{I}_{n}=\sqrt{n}E_{n}[H_{i}h_{i}]$, generalizing the quadratic penalty in Example 2.2. Our first result shows that this minimax design is of the form studied in the Section 3, characterizing the acceptance region induced by this convex penalty.
Theorem 5.1.
The following hold:
  1. (a)
    (Rerandomization). The acceptance criterion $\sup_{\gamma\in B}|\gamma^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]|\leq\epsilon$ $\iff$ $\sqrt{n}E_{n}[H_{i}h_{i}]\in A$ for $A=\epsilon B^{\circ}$ with $B^{\circ}=\{a:\sup_{\gamma\in B}|\gamma^{\prime}a|\leq 1\}\subseteq\mathbb{R}^{d_{h}}$. [^19] Note that for a set $S\subseteq\mathbb{R}^{d}$, we have $\epsilon S=\{\epsilon s:s\in S\}$.
  2. (b)
    (Acceptance Region). $A=\epsilon B^{\circ}$ is symmetric and convex. If $B$ is bounded, then $A$ is closed and has non-empty interior. If $B$ is open, then $A$ is bounded.
  3. (c)
@@
    (Well-specification). If $\gamma_{0}\in B$, then $\operatorname{Var}(R_{A})\leq\epsilon^{2}$.
Part (a) of Theorem 5.1 shows that the rerandomization criterion is of the form studied in Definition 2.1, with acceptance region $A=\epsilon B^{\circ}$. Part (b) shows that $A$ is always symmetric and convex. In particular, the asymptotic distribution of $\widehat{\theta}$ is centered at zero. The set $B^{\circ}$ is known as the absolute polar of $B$, e.g. see Aliprantis and Border, (2006). Part (c) of the theorem shows that if the prior information set $B$ contains the true coefficient $\gamma_{0}$, then $\operatorname{Var}(R_{A})\leq\epsilon^{2}$. Then by independence, the asymptotic variance is within $\epsilon^{2}$ of the optimal partially linear variance. If $\gamma_{0}\not\in B$ (misspecification), then possibly $\operatorname{Var}(R_{A})>\epsilon^{2}$. However, note that misspecification does not affect our inference methods, which allow for general acceptance regions $A$.
Remark 5.2 (Acceptance Probability).
Note that the asymptotic acceptance probability $a(\epsilon)=P(Z_{h}\in\epsilon B^{\circ})$ has $a(\epsilon)\to 0$ as $\epsilon\to 0$ and is monotonically increasing. For $B$ bounded, the theorem shows that $B^{\circ}$ has non-empty interior. In this case, as $\epsilon\to\infty$ we have $\epsilon B^{\circ}\uparrow\mathbb{R}^{d_{h}}$ so $a(\epsilon)\to 1$. This shows that, at least in large samples, we can choose $\epsilon$ to achieve any desired acceptance probability $P(Z_{h}\in A)\in(0,1)$. Under well-specification, any such choice of $\epsilon$ comes with a variance guarantee provided by the theorem.
@@ 5.2 Specifying Prior Information
Without pilot data, we are left to introspection to choose the prior information set $B$. Recall that $\gamma_{0}$ is the coefficient from the partially linear projection $\bar{Y}=\gamma_{0}^{\prime}h+t_{0}(\psi)+e$. Intuitively, $\gamma_{0}$ parameterizes how much we expect the average outcome level to change given a unit change in $h$, holding $\psi$ fixed. If the partially linear model happens to be well-specified, then $\gamma_{0}=\nabla_{b}E[\bar{Y}|\psi,h=b]$. If $t_{0}(\psi)=t^{\prime}\psi$ happens to be linear, then $\bar{Y}=c+\gamma_{0}^{\prime}h+t^{\prime}\psi+e$ and $\gamma_{0}$ is just an OLS coefficient. The following examples provide some reasonable prior information specifications and their associated acceptance regions. These examples rely on a general characterization of acceptance regions in Lemma 5.5 below.
@@
One natural way to specify prior information is to assume $\gamma_{0j}\in[l_{j},u_{j}]$ for each $1\leq j\leq d_{h}$, equivalent to setting $B=\prod_{j=1}^{d_{h}}[l_{j},u_{j}]$. This allows sign constraints, e.g. $0\leq\gamma_{0j}\leq m$ for some $j$ and $-m\leq\gamma_{0j}\leq 0$ for others. Lemma 5.5 below shows that if $B=\prod_{j=1}^{d_{h}}[l_{j},u_{j}]$, then $A=\epsilon B^{\circ}=\{a:|a^{\prime}l+a^{\prime}u|+\sum_{j}|a_{j}|u_{j}-|a_{j}|l_{j}\leq 2\epsilon\}$, where $l=(l_{j})_{j}$ and $u=(u_{j})_{j}$. An example is shown in Figure 1. Note that the acceptance region $A$ is conservative in directions aligned with the prior information set $B=[1,2]\times[1,3/2]$, guarding against covariate imbalances that are aligned with adverse coefficient values $\gamma_{0}\in B$. $A$ is more lenient in directions approximately orthogonal to $B$.
@@
Another natural specification is to guess $\gamma_{0}\approx\bar{\gamma}$, setting $B=\bar{\gamma}+B_{2}(0,m)$, for an uncertainty parameter $m$. By the characterization in Lemma 5.5 below, $A=\epsilon B^{\circ}=\{a:|a^{\prime}\bar{\gamma}|+m|a|_{2}\leq\epsilon\}$. More generally, if $B=\bar{\gamma}+\Sigma B_{2}(0,1)$ for a positive-definite matrix $\Sigma$, the lemma shows that $A=\epsilon B^{\circ}=\{a:|a^{\prime}\bar{\gamma}|+|\Sigma a|_{2}\leq\epsilon\}$. One natural application of this specification is when $B$ is a Wald confidence region constructed using pilot data, as discussed below. An example is shown in Figure 1.
@@
Figure 1: Prior information $B$ and $A=\epsilon B^{\circ}$ for Examples 5.3 and 5.4.
More generally, the following lemma provides a useful characterization of the acceptance region $A=\epsilon B^{\circ}$ from Theorem 5.1 for a large family of prior information set specifications. To state the lemma, recall that $|x|_{p}=(\sum_{j}|x_{j}|^{p})^{1/p}$ for $p\in[1,\infty)$ and $|x|_{\infty}=\max_{j}|x_{j}|$. For $p\in[1,\infty]$, denote $B_{p}(0,1)=\{a:|a|_{p}\leq 1\}$.
@@
For $p\in[1,\infty]$, let $11/p+1/q=$ 1, setting $q=1$ if $p=\infty$ and vice-versa. Suppose $B=x+\Sigma B_{p}(0,1)$, for $x\in\mathbb{R}^{d_{h}}$ and $\Sigma$ invertible. Then $A=\epsilon B^{\circ}=\{a:|a^{\prime}x|+|\Sigma^{\prime}a|_{q}\leq\epsilon\}$.
@@ 5.3 Using Pilot Data
Next, we discuss an alternative strategy that uses pilot data to specify the set $B$. Suppose we have access to $\mathcal{D}_{pilot}\perp\!\!\!\perp(W_{1:n},D_{1:n})$ of size $m$. Suppose $\sqrt{m}(\widehat{\gamma}_{pilot}-\gamma_{0})\approx\mathcal{N}(0,\widehat{\Sigma}_{pilot})$ for some estimator $\widehat{\gamma}_{pilot}$, discussed below. Consider forming the Wald region $\widehat{B}_{pilot}=\{\gamma:m(\widehat{\gamma}_{pilot}-\gamma)^{\prime}\widehat{\Sigma}_{pilot}^{-1}(\widehat{\gamma}_{pilot}-\gamma)\leq c_{\alpha}\}$ using critical value $P(\chi^{2}_{d_{h}}\leq c_{\alpha})=1-\alpha$ for $\alpha\in(0,1)$. Equivalently, one can write the Wald region as
@@
Using $\widehat{B}_{pilot}$ as a prior information set, by Example 5.4 we have acceptance region
$\widehat{A}_{pilot}=\epsilon\widehat{B}_{pilot}^{\circ}=\{a:|a^{\prime}\widehat{\gamma}_{pilot}|+m^{-1/2}c_{\alpha}^{1/2}|\widehat{\Sigma}^{1/2}a|_{2}\leq\epsilon\}.$ (5.3)
Note that the acceptance region $\widehat{A}_{pilot}$ grows with the pilot size $m$. This reflects smaller uncertainty about the true parameter $\gamma_{0}$, and thus less adversarial worst case imbalance $\sup_{\gamma\in\widehat{B}_{pilot}}|\gamma^{\prime}E_{n}[H_{i}h_{i}]|$. Conversely, $\widehat{A}_{pilot}$ shrinks as the confidence parameter $\alpha$ and the scale of the variance estimate $\widehat{\Sigma}_{pilot}$ increases, reflecting greater uncertainty and a more conservative approach to covariate balances. Our next result shows that rerandomization with acceptance region $\widehat{A}_{pilot}$ controls the variance of the imbalance $R_{A}=\gamma_{0}^{\prime}Z|Z\in\widehat{A}_{pilot}$ with high probability marginally over the realizations of the pilot data. The result is an immediate consequence of Theorem 3.4 and Theorem 5.1.
@@
Suppose $P(\gamma_{0}\in\widehat{B}_{pilot})\geq 1-\alpha$, for $\mathcal{D}_{pilot}\perp\!\!\!\perp(W_{1:n},D_{1:n})$. Let $D_{1:n}$ as in Definition 2.1 with $A=\widehat{A}_{pilot}=\epsilon\widehat{B}_{pilot}^{\circ}$, then if Assumptions 3.1, 3.2 hold, then $\sqrt{n}(\widehat{\theta}-\theta_{n})|\mathcal{D}_{pilot}\Rightarrow v_{D}^{-1}\mathcal{N}(0,\operatorname{Var}(e))+R_{A}$ and $\operatorname{Var}(R_{A}|\mathcal{D}_{pilot})\leq\epsilon^{2}$ with probability $\geq 1-\alpha$.
@@
General Parameters. For completeness, we extend the preceding work to general parameters $\theta_{n}$ as in Definition 2.5. Let $\Pi\delta(W,\theta_{0})$ be the assignment influence function. As in Equation 3.4, consider the partially linear decomposition
$\Pi\delta(W,\theta_{0})=\Gamma_{0}^{\prime}h+t_{0}(\psi)+e,\quad\;E[e|\psi]=0,\;E[eh]=0.$
@@
Let $D_{1:n}$ as in Definition 2.1 with $A=\cap_{j=1}^{d_{\theta}}\epsilon B^{\circ}_{j}$. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R_{A}$, as defined in Theorem 3.4. If $\Gamma_{0}^{j}\in B_{j}$ $\forall j$, then $\max_{j=1}^{d_{\theta}}\operatorname{Var}((R_{A})_{jj})\leq\epsilon^{2}$.
@@
6 Linear Adjustment
In this section, we study optimal linearly adjusted GMM estimation under stratified rerandomization. We show that this can be used to completely remove the impact of the acceptance region and imbalance term $R_{A}$ to first order, restoring asymptotic normality. This allows for standard t-statistic and Wald-test based inference on the parameters $\theta_{n}$ and $\theta_{0}$, provided in Section 7 below. Let $w$ denote the covariates used for ex-post adjustment and suppose $E[|w|_{2}^{2}]<\infty$.

### v2

@@ 5.1 Minimax Rerandomization
Since $\gamma_{0}$ is unknown at design-time, we instead take a minimax approach that incorporates prior information about the coefficient $\gamma_{0}$. For information set $B\subseteq\mathbb{R}^{d_{h}}$ specified by the researcher, consider rerandomizing until the worst case in-sample correlation between treatments and covariates is small enough,
$\sup_{\gamma\in B}|\gamma^{\prime}\mathcal{I}_{n}|=\sup_{\gamma\in B}\left|\frac{\operatorname{Cov}_{n}(\gamma^{\prime}h_{i},D_{i})}{\operatorname{Var}_{n}(D_{i})}\right|\leq\epsilon.$ (5.1)
Equivalently, we rerandomize until $p_{B}(\mathcal{I}_{n})\leq\epsilon$ for the convex penalty function $p_{B}(x)=\sup_{\gamma\in B}|\gamma^{\prime}x|$. This significantly generalizes the commonly used quadratic penalty function $p(x)=x^{\prime}\Sigma x$ (e.g. Schindl and Branson, (2024)). Our next result shows that Equation 5.1 is a linear rerandomization, characterizing the implicit acceptance region $A$.
Proposition 5.1 (Acceptance Region).
The criterion $p_{B}(\mathcal{I}_{n})\leq\epsilon$ $\iff$ $\mathcal{I}_{n}\in A_{0}$ for $A_{0}=\epsilon B^{\circ}$ with $B^{\circ}=\{x:\sup_{\gamma\in B}|\gamma^{\prime}x|\leq 1\}\subseteq\mathbb{R}^{d_{h}}$, the absolute polar of $B$. The set $A_{0}$ is symmetric and convex. If $B$ is bounded, $A_{0}$ is closed and has non-empty interior. [^23] Also if $\operatorname{int}B\not=\emptyset$ then $A_{0}$ is bounded. See Aliprantis and Border, (2006) for more on polar sets.
Note that since $A_{0}$ is symmetric, the discussion after Theorem 3.5 implies that the asymptotic distribution of $\widehat{\theta}$ under the design in Equation 5.1 is centered at zero. We let $B$ be totally bounded in what follows. The proposition shows that in this case $A_{0}$ has non-empty interior, satisfying the conditions of Assumption 3.1.
We showed above that $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow L_{0}$ for $L_{0}=\mathcal{N}(0,V(\gamma_{0}))+\gamma_{0}^{\prime}Z_{hA}$. Since $\gamma_{0}$ is unknown at design time, define a family $\{L_{\gamma,A}:\gamma\in\mathbb{R}^{d_{h}},A\subseteq\mathbb{R}^{d_{h}}\}$ of possible limiting distributions of $\widehat{\theta}$, with each $L_{\gamma,A}=\mathcal{N}(0,V(\gamma))+\gamma^{\prime}Z_{hA}$ a sum of independent RV’s as above. For any distribution in this family, the conditional asymptotic bias of $\widehat{\theta}$ given realized covariate imbalances $Z_{hA}$ is $\operatorname{bias}(L_{\gamma,A}|Z_{hA})\equiv E[L_{\gamma,A}|Z_{hA}]$. Our main result shows that the polar acceptance region $A_{0}=\epsilon B^{\circ}$ maximizes rerandomization acceptance probability $P(Z_{h}\in A)$, subject to a strict constraint on asymptotic conditional bias, uniformly over all limiting distributions consistent with the prior information $\gamma_{0}\in B$.
Theorem 5.2 (Minimax).
The acceptance region $A_{0}=\epsilon B^{\circ}$ solves [^24] Implicitly, we maximize only over Borel-measurable sets $A\in\mathcal{B}(\mathbb{R}^{d_{h}})$. The solution $A_{0}$ is unique up to the equivalence class $\{A\in\mathcal{B}(\mathbb{R}^{d_{h}}):\operatorname{Leb}(A\triangle A_{0})=0\}$, where $\triangle$ denotes symmetric difference.
$A_{0}=\operatorname*{argmax}_{A\subseteq\mathbb{R}^{d_{h}}}P(Z_{h}\in A)\quad\text{s.t.}\quad\,\sup_{\gamma\in B}|\operatorname{bias}(L_{\gamma,A}|Z_{hA})|\leq\epsilon.$ (5.2)
@@
In particular, if $\gamma_{0}\in B$ (well-specification) then $|\operatorname{bias}(L_{0}|Z_{hA_{0}})|\leq\epsilon$ and $\operatorname{Var}(L_{0})\leq V_{a}+\epsilon^{2}$, where $V_{a}$ is the partially linear variance in Equation 3.6.
The expected number of independent draws of $Z_{h}$ until $Z_{h}\in A$ is $P(Z_{h}\in A)^{-1}$. Then, equivalently, Equation 5.2 shows that the acceptance region $A_{0}$ minimizes the (asymptotic) computational cost of rerandomization, subject to a strict guarantee on statistical performance. From the final statement of the theorem, if the prior information set $B$ is well-specified ($\gamma_{0}\in B$), setting $A_{0}=\epsilon B^{\circ}$ bounds the magnitude of the conditional asymptotic bias $E[L_{0}|Z_{hA_{0}}]$ of the GMM estimator $\widehat{\theta}$ below $\epsilon$. In particular, this implies that the variance $\operatorname{Var}(L_{0})$ of the asymptotic distribution $\sqrt{n}(\widehat{\theta}-\theta_{n})\Rightarrow L_{0}=\mathcal{N}(0,V_{a})+\gamma_{0}^{\prime}Z_{hA_{0}}$ is within $\epsilon^{2}$ of the optimal partially linear variance $V_{a}$ in Equation 3.7.
Remark 5.3 (Integral Probability Metric).
Before continuing, we briefly note an interesting interpretation of the design in Equation 5.1. For distributions $P,Q$ and a function class $\mathcal{F}$, the integral probability metric is $\rho(P,Q;\mathcal{F})\equiv\sup_{f\in\mathcal{F}}|E_{P}[f(X)]-E_{Q}[f(X)]|$. [^25] The pseudometric $\rho$ is also referred to as the maximum mean discrepancy. This is a commonly used statistic in two-sample testing, see e.g. Gretton et al., (2008). Let $\mathcal{F}_{B}=\{\gamma^{\prime}h:\gamma\in B\}$ and define empirical distributions $\widehat{P}_{d}=(h_{i}|D_{i}=d)$ for $d=0,1$. We have
$\sup_{\gamma\in B}|\gamma^{\prime}\mathcal{I}_{n}|\leq\epsilon\iff\sqrt{n}\rho(\widehat{P}_{1},\widehat{P}_{0};\mathcal{F}_{B})\leq\epsilon.$
The minimax design rerandomizes until covariates $h$ are balanced according to $\rho(\widehat{P}_{1},\widehat{P}_{0};\mathcal{F}_{B})$, a distance between covariate distributions that is only sensitive to the projections $\gamma^{\prime}h$ that actually matter for estimating $\theta_{n}=\operatorname{SATE}$. By doing so, we maximize acceptance probability subject to the statistical guarantee in Theorem 5.2.
@@ 5.2 Specifying Prior Information
Without pilot data, we must use introspection to choose the prior information set $B$ containing $\gamma_{0}$. Intuitively, $\gamma_{0}$ parameterizes how much the average outcome level $\bar{Y}$ can change for a unit change in $h$, holding $\psi$ fixed. If $t_{0}(\psi)=t^{\prime}\psi$ happens to be linear, then $\bar{Y}=c+\gamma_{0}^{\prime}h+t^{\prime}\psi+e$ and $\gamma_{0}$ is an OLS coefficient. The following examples provide some reasonable prior information specifications and their associated acceptance regions, using a general characterization of acceptance regions $A_{0}=\epsilon B^{\circ}$ in Lemma 5.6 below.
@@
One specification is to assume $\gamma_{0j}\in[l_{j},u_{j}]$ for each $1\leq j\leq d_{h}$, setting $B=\prod_{j=1}^{d_{h}}[l_{j},u_{j}]$. This allows for sign constraints, e.g. $0\leq\gamma_{0j}\leq m$ for some $j$ and $-m\leq\gamma_{0j}\leq 0$ for others. Lemma 5.6 below shows that if $B=\prod_{j=1}^{d_{h}}[l_{j},u_{j}]$, then $A_{0}=\epsilon B^{\circ}=\{x:|x^{\prime}l+x^{\prime}u|+\sum_{j}|x_{j}|u_{j}-|x_{j}|l_{j}\leq 2\epsilon\}$, where $l=(l_{j})_{j}$ and $u=(u_{j})_{j}$. An example is shown in Figure 2. Note that the region $A_{0}$ is conservative in the direction of the set $B=[1,2]\times[1,3/2]$, rejecting covariate imbalances that are too closely aligned with adverse coefficient values $\gamma_{0}\in B$. The region $A_{0}$ is more lenient in directions approximately orthogonal to $B$.
@@
Another natural specification is to set $B=\bar{\gamma}+B_{2}(0,m)$, for an uncertainty parameter $m$ and guess $\gamma_{0}\approx\bar{\gamma}$. By Lemma 5.6, $A_{0}=\{x:|x^{\prime}\bar{\gamma}|+m|x|_{2}\leq\epsilon\}$. More generally, if $B=\bar{\gamma}+\Sigma B_{2}(0,1)$ for a positive-definite matrix $\Sigma$, then $A_{0}=\{x:|x^{\prime}\bar{\gamma}|+|\Sigma x|_{2}\leq\epsilon\}$. One application of this specification is when $B$ is a Wald confidence region constructed using pilot data, as discussed below.
@@
Figure 2: Prior information $B$ and $A_{0}=\epsilon B^{\circ}$ for Examples 5.4 and 5.5.
The following lemma provides a useful characterization of the acceptance region $A_{0}=\epsilon B^{\circ}$ from Theorem 5.2 for a large family of prior information set specifications. To state the lemma, recall that $|x|_{p}=(\sum_{j}|x_{j}|^{p})^{1/p}$ for $p\in[1,\infty)$ and $|x|_{\infty}=\max_{j}|x_{j}|$. For $p\in[1,\infty]$, denote $B_{p}(0,1)=\{x:|x|_{p}\leq 1\}$.
@@
For $p\in[1,\infty]$, let $11/p+1/q=$ 1, setting $q=1$ if $p=\infty$ and vice-versa. Suppose $B=b+\Sigma B_{p}(0,1)$, for $b\in\mathbb{R}^{d_{h}}$ and $\Sigma$ invertible. Then $A_{0}=\{x:|x^{\prime}b|+|\Sigma^{\prime}x|_{q}\leq\epsilon\}$.
@@ 5.3 Using Pilot Data
Next, we discuss an alternative strategy that uses pilot data to specify the set $B$. Suppose we have access to $\mathcal{D}_{pilot}\perp\!\!\!\perp(W_{1:n},D_{1:n})$ of size $m$. Suppose $\sqrt{m}(\widehat{\gamma}_{pilot}-\gamma_{0})\approx\mathcal{N}(0,\widehat{\Sigma}_{pilot})$ for some pilot estimator $\widehat{\gamma}_{pilot}$, discussed below. Consider forming the Wald region $\widehat{B}_{pilot}=\{\gamma:m(\widehat{\gamma}_{pilot}-\gamma)^{\prime}\widehat{\Sigma}_{pilot}^{-1}(\widehat{\gamma}_{pilot}-\gamma)\leq c_{\alpha}\}$ using critical value $P(\chi^{2}_{d_{h}}\leq c_{\alpha})=1-\alpha$ for $\alpha\in(0,1)$. Equivalently, one can write this Wald region as
@@
By Example 5.5, using $\widehat{B}_{pilot}$ as a prior information set gives acceptance region
$\widehat{A}_{pilot}=\epsilon\widehat{B}_{pilot}^{\circ}=\{x:|x^{\prime}\widehat{\gamma}_{pilot}|+\frac{c_{\alpha}^{1/2}|\widehat{\Sigma}^{1/2}x|_{2}}{m^{1/2}}\leq\epsilon\}.$ (5.4)
Note that the acceptance region $\widehat{A}_{pilot}$ expands as the pilot size $m$ is larger. This reflects smaller uncertainty about the true parameter $\gamma_{0}$, and thus less adversarial worst case imbalance $\sup_{\gamma\in\widehat{B}_{pilot}}|\gamma^{\prime}\mathcal{I}_{n}|$. Conversely, $\widehat{A}_{pilot}$ shrinks as the confidence parameter $\alpha$ and the scale of the variance estimate $\widehat{\Sigma}_{pilot}$ increases, reflecting greater uncertainty and a more conservative approach to covariate balances. Our next result shows that rerandomization with acceptance region $\widehat{A}_{pilot}$ controls the variance of the residual imbalance $R_{A}=\gamma_{0}^{\prime}Z|Z\in\widehat{A}_{pilot}$ with high probability, marginally over the realizations of the pilot data. The result is an immediate consequence of Theorem 3.5 and Theorem 5.2.
@@
Suppose $P(\gamma_{0}\in\widehat{B}_{pilot})\geq 1-\alpha$, for $\mathcal{D}_{pilot}\perp\!\!\!\perp(W_{1:n},D_{1:n})$. Let $D_{1:n}$ as in Definition 2.1 with $A=\widehat{A}_{pilot}=\epsilon\widehat{B}_{pilot}^{\circ}$. If Assumptions 3.1, 3.2 hold, then $\sqrt{n}(\widehat{\theta}-\theta_{n})|\mathcal{D}_{pilot}\Rightarrow v_{D}^{-1}\mathcal{N}(0,\operatorname{Var}(e))+R_{A}$, where $\operatorname{Var}(R_{A}|\mathcal{D}_{pilot})\leq\epsilon^{2}$ with probability $\geq 1-\alpha$.
@@
General Parameters. For completeness, we extend the preceding work to general parameters $\theta_{n}$ as in Definition 2.5. Let $\Pi a(W,\theta_{0})$ be the assignment influence function. As in Equation 3.7, consider the partially linear decomposition
$\Pi a(W,\theta_{0})=\gamma_{0}^{\prime}h+t_{0}(\psi)+e,\quad\;E[e|\psi]=0,\;E[eh]=0.$
@@
Let $D_{1:n}$ as in Definition 2.1 with $A=\cap_{j=1}^{d_{\theta}}\epsilon B^{\circ}_{j}$. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R_{A}$, as defined in Theorem 3.5. If $\gamma_{0}^{j}\in B_{j}$ $\forall j$, then $\max_{j=1}^{d_{\theta}}\operatorname{Var}((R_{A})_{jj})\leq\epsilon^{2}$.
@@
6 Restoring Normality
In this section, we study optimal linearly adjusted GMM estimation under stratified rerandomization. We show that, to first order, optimal ex-post linear adjustment completely removes the impact of the acceptance region $A$ and imbalance term $R_{A}$, restoring asymptotic normality. This enables standard t-statistic and Wald-test based inference on the parameters $\theta_{n}$ and $\theta_{0}$, provided in Section 7 below.
Let $w$ denote the covariates used for ex-post adjustment and suppose $E[|w|_{2}^{2}]<\infty$.